# APTQ: Attention-aware Post-Training Mixed-Precision Quantization for Large Language Models

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: DAC 2024: 61st IEEE/ACM Design Automation Conference.; June 23-27; San Francisco, CAisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generation
@@ 1. Introduction
Large Language Models (LLMs), such as ChatGPT (Ouyang et al., 2022), LLaMA (Touvron et al., 2023), etc., exhibit impressive performance across various tasks. However, deploying these models on edge devices is challenging due to their exorbitant computational demands and memory footprints. Existing model compression solutions such as pruning (Carreira-Perpinán and Idelbayev, 2018) and neural architecture search (Chen et al., 2021) often require model retraining, which is extremely time-consuming and expensive for billion-parameter models. Recently, post-training quantization (PTQ) methods, such as GPTQ (Frantar et al., 2023), have been proposed and achieved relatively high accuracy without retraining. However, GPTQ only considers the weight quantization strategy in the scope of a single layer as an optimization problem to minimize $\bm{W}\bm{X}-\bm{\hat{W}}\bm{X}||_{2}^{2}$, with $\bm{W}$, $\bm{\hat{W}}$ and $\bm{X}$ representing float weights, quantized weights and inputs, respectively. This simplification fails to consider the complex and nonlinear effects such as softmax in the attention computation, and leads to a sub-optimal solution.
@@ 2. Related Work
To deploy large models on edge devices, quantization is a versatile technique for reducing model size and computation. Quantization-Aware Training (QAT) is known to be effective by integrating the quantization process into the training process. A representative work is LLM-QAT (Liu et al., 2023b), which proposes data-free distillation. However, this method introduces new trainable parameters, necessitates high-end GPU computational resources, and incurs a large time consumption. In contrast, Post-Training Quantization (PTQ) employs moderate resources to quantize pre-trained models without model retraining. SqueezeLLM (Kim et al., 2023), compress most weights to 4 bits but maintain outlier weights at 16 bits, which complicates the inference process with both 4-bit and 16-bit inference.
SmoothQuant (Xiao et al., 2023) introduces a per-channel scaling transformation that effectively smooths the magnitudes to address the challenge of quantizing activations. GPTQ (Frantar et al., 2023) and OBQ (Frantar and et al., 2022) introduce an innovative weight quantization method based on approximate second-order information, ensuring high accuracy and efficiency in the quantization process. Our work shares the same ethos as GPTQ but additionally considers the softmax and matmul operations within the attention computation to formulate the quantization problem, resulting in improved accuracy.
@@ 3. Algorithm
This section starts with the preliminaries to outline the evolution of quantization techniques from optimal brain quantization (OBQ) (Frantar and et al., 2022) to our proposed Hessian-attention-based quantization. We then propose an Attention-aware Post-Training Mixed-Precision Quantization, APTQ, to further compress the LLMs.
@@ 3.1. Preliminaries
Optimal Brain Quantization (OBQ). Optimal Brain Quantization (OBQ) (Frantar and et al., 2022) is an innovative method that minimizes quantization errors by treating each neural network weight independently. The core of OBQ lies in iteratively quantizing each weight and adjusting the remaining unquantized weights to compensate for the quantization-induced errors. This approach is mathematically articulated as follows:
@@ 3.2. Hessian-Attention-based Quantization
[…]
@@
Derivatives for Different Quantization Layers. The current problem is transformed into finding the partial derivative of F(W^)F($\hat${W}) with respect to the weights $\hat{W}$. The F(W^)F($\hat${W}) function is different for the Feed-Forward layers and Attention layers. In the Feed-Forward layer, the main structure is a linear fully connected layer. The Hessian matrix is easily computed as $H_{F}=2X_{F}X_{F}^{T}$, corresponding to the Hessian matrix form in the GPTQ method.
@@
Here, $W_{h}$ represents the weight matrix in the $n$-th attention head, and NN and $\mathbb{P}_{h}$ are given by:
@@
Here,E{E} represents the quantization error, $w_{q}$ refers to the quantized weights of the current group. $\delta_{F}$ refers to the corresponding optimal updates for the remaining float weights (not yet quantized weights of the current layer). This principle is uniformly applicable to the quantization of $Q$ (query), $K$ (key), $V$ (value), and $O$ (output) weight matrices in attention mechanisms. By synthesizing these elements, we can effectively compute the second-order Hessian information relevant to the weights within the attention layers. This advanced computation aids in the update and optimization of weights, targeting the minimization of the original squared error as defined in equation (5). This approach facilitates the realization of quantized models with robust performance across different components of the attention mechanism. The comprehensive algorithm is detailed in Algorithm Box 1.
@@
Input: Pre-trained model weights $W$, blocksize $B$, Hessian matrix $H$, quantization function quant, Layer names layerNamelayerName, Ratio of 4-bit in 2/4 mixed-precision $R$.
@@
$3.$ Hessian-Trace-based Mixed-Precision Quantization
@@
(18) $\text{average bits}=4\times R+2\times(1-R),$
@@
Method Avg bit $\downarrow$ C4 $\downarrow$ WikiText-2 $\downarrow$
@@ 4.3. Evaluation of Zero-shot performance
As depicted in Table 2, we benchmark the APTQ framework against current SOTA PTQ methodologies applied to the LLaMa-7B model. Our findings illustrate that APTQ, when configured to 3.8 bits, sustains a remarkably minimal deviation in accuracy, with a diminutive average accuracy drop of only 0.33 points from the full-precision model. Even when the APTQ is optimized down to an average of 3.6 or 3.5 bits, it still consistently outperforms the majority of 4-bit PTQ models. These findings demonstrate that APTQ excels in zero-shot tasks with minimal bit usage, highlighting its effectiveness in deploying large-scale language models in environments with limited computational resources. This underscores APTQ’s advantage in resource-efficient performance.
$4.$ Ablation Study

### v2

@@ Abstract.
Equal contribution.
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: 61st ACM/IEEE Design Automation Conference; June 23–27, 2024; San Francisco, CA, USAbooktitle: 61st ACM/IEEE Design Automation Conference (DAC ’24), June 23–27, 2024, San Francisco, CA, USAdoi: 10.1145/3649329.3658498isbn: 979-8-4007-0601-1/24/06ccs: Computing methodologies Natural language generation
@@ 1. Introduction
Large Language Models (LLMs), such as ChatGPT (Ouyang et al., 2022), OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023), etc., exhibit impressive performance across various tasks. However, deploying these models on edge devices is challenging due to their exorbitant computational demands and memory footprints. Existing model compression solutions such as pruning (Carreira-Perpinán and Idelbayev, 2018) and neural architecture search (Chen et al., 2021) often require model retraining, which is extremely time-consuming and expensive for billion-parameter models. Recently, post-training quantization (PTQ) methods, such as GPTQ (Frantar et al., 2023), have been proposed and achieved relatively high accuracy without retraining. However, GPTQ only considers the weight quantization strategy in the scope of a single layer as an optimization problem to minimize $||\bm{W}\bm{X}-\bm{\hat{W}}\bm{X}||_{2}^{2}$, with $\bm{W}$, $\bm{\hat{W}}$ and $\bm{X}$ representing float weights, quantized weights and inputs, respectively. This simplification fails to consider the complex and nonlinear effects such as softmax in the attention computation, and leads to a sub-optimal solution.
@@ 2. Related Work
To deploy large models on edge devices, quantization is a versatile technique for reducing model size and computation. Quantization-Aware Training (QAT) is known to be effective by integrating the quantization process into the training process. A representative work is LLM-QAT (Liu et al., 2023b), which proposes data-free distillation. However, this method introduces new trainable parameters, necessitates high-end GPU computational resources, and incurs a large time consumption. In contrast, Post-Training Quantization (PTQ) employs moderate resources to quantize pre-trained models without model retraining. Recent work, such as SpQR (Dettmers et al., 2023) and SqueezeLLM (Kim et al., 2023), compress most weights to 4 bits but maintain outlier weights at 16 bits, which complicates the inference process with both 4-bit and 16-bit inference.
SmoothQuant (Xiao et al., 2023) introduces a per-channel scaling transformation that effectively smooths the magnitudes to address the challenge of quantizing activations. GPTQ (Frantar et al., 2023) and OBQ (Frantar and Alistarh, 2022) introduce an innovative weight quantization method based on approximate second-order information, ensuring high accuracy and efficiency in the quantization process. Our work shares the same ethos as GPTQ but additionally considers the softmax and matmul operations within the attention computation to formulate the quantization problem, resulting in improved accuracy.
@@ 3. Algorithm
This section starts with the preliminaries to outline the evolution of quantization techniques from optimal brain quantization (OBQ) (Frantar and Alistarh, 2022) to our proposed Hessian-attention-based quantization. We then propose an Attention-aware Post-Training Mixed-Precision Quantization, APTQ, to further compress the LLMs.
@@ 3.1. Preliminaries
Optimal Brain Quantization (OBQ). Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022) is an innovative method that minimizes quantization errors by treating each neural network weight independently. The core of OBQ lies in iteratively quantizing each weight and adjusting the remaining unquantized weights to compensate for the quantization-induced errors. This approach is mathematically articulated as follows:
@@ 3.2. Hessian-Attention-based Quantization
(6) $H_{\hat{W}}=2\cdot\left(F^{\prime}(\hat{W})\cdot F^{\prime}(\hat{W})^{T}+[F(W)-F(\hat{W})]\cdot F^{\prime\prime}(\hat{W})\right).$
@@
Derivatives for Different Quantization Layers. The current problem is transformed into finding the partial derivative of $F(\hat{W})$ with respect to the weights $\hat{W}$. The $F(\hat{W})$ function is different for the Feed-Forward layers and Attention layers. In the Feed-Forward layer, the main structure is a linear fully connected layer. The Hessian matrix is easily computed as $H_{F}=2X_{F}X_{F}^{T}$, corresponding to the Hessian matrix form in the GPTQ method.
@@
Here, $W_{h}$ represents the weight matrix in the $n$-th attention head, and $N$ and $\mathbb{P}_{h}$ are given by:
@@
Here,${E}$ represents the quantization error, $w_{q}$ refers to the quantized weights of the current group. $\delta_{F}$ refers to the corresponding optimal updates for the remaining float weights (not yet quantized weights of the current layer). This principle is uniformly applicable to the quantization of $Q$ (query), $K$ (key), $V$ (value), and $O$ (output) weight matrices in attention mechanisms. By synthesizing these elements, we can effectively compute the second-order Hessian information relevant to the weights within the attention layers. This advanced computation aids in the update and optimization of weights, targeting the minimization of the original squared error as defined in equation (5). This approach facilitates the realization of quantized models with robust performance across different components of the attention mechanism. The comprehensive algorithm is detailed in Algorithm Box 1.
@@
Input: Pre-trained model weights $W$, blocksize $B$, Hessian matrix $H$, quantization function quant, Layer names $layerName$, Ratio of 4-bit in 2/4 mixed-precision $R$.
@@
3.3. Hessian-Trace-based Mixed-Precision Quantization
@@
(18) average $\text{average bits}=4\times R+2\times(1-R),$
@@ 4.1. Experiment Setup
Method Avg bit C4 $\downarrow$ WikiText-2 $\downarrow$
@@ 4.3. Evaluation of Zero-shot performance
As depicted in Table 2, we benchmark the APTQ framework against current SOTA PTQ methodologies applied to the LLaMa-7B model. Our findings illustrate that APTQ, when configured to 3.8 bits, sustains a remarkably minimal deviation in accuracy, with a diminutive average accuracy drop of only 0.32 points from the full-precision model. Even when the APTQ is optimized down to an average of 3.6 or 3.5 bits, it still consistently outperforms the majority of 4-bit PTQ models. These findings demonstrate that APTQ excels in zero-shot tasks with minimal bit usage, highlighting its effectiveness in deploying large-scale language models in environments with limited computational resources. This underscores APTQ’s advantage in resource-efficient performance.
4.4. Ablation Study